# Bridging Today and the Future of Humanity: AI Safety in 2024 and Beyond

### v3

@@ 5.3 AI Privacy
Privacy leakage in AI models refers to the unintended exposure of sensitive information that the models may have learned during training. Adversaries can exploit model weights or gradients to infer sensitive information in the training data, e.g., data reconstruction attacks, membership inference attacks, etc [299, 90, 291]. Additionally, AI models are often trained on massive datasets that may contain personally identifiable information (PII) or other confidential data. As a result, these models might memorize specific details of personal information, which might be reproduced inadvertently during interactions with users. Furthermore, malicious users can craft adversarial prompts to extract sensitive information from AI models, posing risks to data privacy and security, e.g., jailbreaking [121, 38].
@@
Addressing privacy leakage in AI models involves implementing methods such as differential privacy [87, 242, 296, 294], federated learning [300, 292, 81, 290, 162], and privacy-preserving cryptographic protocols (e.g., homomorphic encryption [115, 184, 99]) for training AI models, ensuring that sensitive data cannot be extracted from weights or model outputs. Researchers also leverage differential privacy in in-context learning and finetuning [284, 293, 227], or design privacy-preserving prompts for querying LLMs [86, 55], to maintain data confidentiality during interactions between users and AI models. Other practices for ensuring data privacy include conducting rigorous data filtering before training, monitoring model inputs and outputs (i.e., moderation in §5.2), etc, to ensure that the model outputs would not reveal sensitive information [137, 116, 83, 94] .
@@
When reflecting on the current practices of data privacy, a crucial aspect is the balance between privacy and utility. While ensuring data privacy is essential, it often comes at the expense of functionality, utility, and user experiences. In practice, implementing privacy measures in AI applications may fail to prevent malicious users while inadvertently degrading the experiences of benign users. Moreover, with advancements in technologies, breaching privacy is becoming easier, while ensuring robust privacy protections is growing more complex and expensive. This highlights a deeper issue: those intent on exploiting vulnerabilities of AI models can often find new, sophisticated methods to bypass safeguards, whereas being overly cautious may impact regular users negatively.
@@ 5.4 AI Security and Defense Methods
Defense methods for ensuring robust AI models and secure model outputs typically involve the following strategies to: (1) aligning models with human values and ethical norms through Supervised Fine-Tuning (SFT) [298] and Reinforcement Learning from Human Feedback (RLHF) [161] to ensure the models to follow user instructions safely and responsibly; (2) leveraging built-in safeguards for content generation, e.g., exploring the decoding stage [91, 54], implementing reward-based mechanisms [147, 192], and leveraging hidden states [29, 54], to prevent the generation of unsafe content; (3) constructing datasets with safety-enhancing features, e.g., secure instructions and adversarial samples, and training models with such data to enhance their robustness against harmful inputs [271, 69, 137, 10, 172]; (4) employing prompt engineering methods, combined with chain-of-thought reasoning or multi-agent systems, to help models understand user inputs better and respond appropriately to user inputs while rejecting inappropriate outputs [180, 265, 254, 288]; (5) training models to detect unsafe content and incorporating moderation pipelines to help block unsafe model inputs and prevent the generation of inappropriate outputs (see §5.2); (6) employing certified robustness approaches that provide formal guarantees on the model’s robustness against adversarial perturbations [77, 42]; and (7) utilizing differential privacy [1, 96, 214], homomorphic encryption [223, 99], and secure MPC [46, 109], to ensure that sensitive data are protected during model training, etc.
@@ 5.5 AI Serving Security
AI serving security involves strategies to ensure secure deployment, operation, and maintenance of AI systems when providing services to users [166, 266]. It focuses on protecting models and interactions between users and models from various threats, such as such as supply chain vulnerabilities [168], model theft [165], model denial of service [164], and insecure plugins [163].
Supply chain vulnerabilities exist in software components, models, and training data provided by third party providers, or user prompts collected with the supply chain [168]. Such vulnerabilities include using outdated models, deprecated third-party packages, or improper handling of training data, leading to data breaches, malicious injections, or even system failures [168]. Hackers may inject poisoned data through supply chain, introducing backdoors and biases during pretraining or fine-tuning [169]. Malicious users may submit malicious prompts to bypass system controls, inducing unauthorized actions or unsafe model outputs [167].
@@
Distributed denial of service (DDoS) attacks compromises AI services, especially when models are integrated into widely accessible applications [164]. Attackers can craft prompts that exploit recursive behaviors in AI models, which consumes excessive computational resources [195]. For example, an abnormal traffic pattern of a DDoS attack led to sporadic outages for OpenAI, affecting ChatGPT and developer tools for hours in November 2023 [170].
Plugins enhance model functionality by enabling interactions with external software, databases, web tools, or APIs [163]. However, they may introduce vulnerabilities when execution controls are inadequate. Adversaries may exploit insecure plugins with adversarial prompts to perform unauthorized actions, such as data exfiltration, remote code execution, and privilege escalation [163]. Additionally, insufficient input validation may cause the unintended disclosure of sensitive information [196]. Furthermore, without isolated environments, unauthorized access to plugins could potentially modify system-level resources, which, in the worst cases, might cause the AI system to crash [196, 197].
@@ 5.6 AI Governance
AI governance is a framework of rules, practices, policies, and technological tools that ensure AI systems are built, developed, and used in a safe and responsible manner, aligning with social values and fulfilling legal standards [136, 21]. [219] defines 13 AI governance components in a continuous loop in AI life-cycle, including: 1) establishing accountability for AI; 2) assessing regulatory risks; 3) gathering inventory of use cases; 4) increasing value of underlying data; 5) assessing fairness and accessibility; 6) improving reliability and safety; 7) heightening transparency and explainability; 8) implementing accountability with human-in-the-loop; 9) supporting privacy and retention; 10) improving security; 11) implementing AI model lifecycle and registry; 12) managing risk; and 13) realizing AI value.
@@
Among these directions, items 1) and 8) are related to accountability; items 2), 6), 9), 10), and 12) are related to jailbreaking and red teaming (see §5.1), content moderation and customized moderation pipeline (see §5.2), AI privacy (see §5.3), AI security and defense methods (see §5.4), and AI serving security (see §5.5); item 3) is related to business and is out of the scope of this subsection (see [219] for details on practical AI governance use cases); item 4) is related to data management such as training data processing [137], access control [6, 16, 160, 208], and data regulations [14, 247]; item 5) is related to customized AI workflows and moderation pipelines (see §5.2) and access control in AI systems [6, 101]; item 7) is related to AI transparency; and items 11) and 13) are related to customizing AI workflow to enhance functionality and safety in AI systems (see §5.2). Without loss of generality, this subsection discusses topics that are not mentioned in the previous subsections, including i) accountability, ii) regulations, iii) access control, and iv) transparency, interpretability and explainability.
@@
Accountability. Accountability refers to being responsible for actions and impacts of AI systems on individuals and society [191, 142]. It ensures responsibility and auditability, enabling redress and minimizing negative impact of AI systems [142]. [153] defines accountability as a relation of answerability with three conditions: authority recognition, interrogation, and limitation of power. It outlines a framework with seven features, including context, range, agent, forum, standards, process, and implications, as well as four key goals, including compliance, reporting, oversight, and enforcement. These goals are often complementary, while policy-makers tend to focus on some goals over others, depending on specific objectives of AI governance.
@@
Accountability can be explored at different levels within the AI lifecycle, including data, model, and developers, and across various sectors such as government, private industry, and academia [190]. Typical approaches for enhancing accountability include algorithmic assessments, auditing, and data provenance techniques [190]. Watermarking methods further support accountability by embedding traceable identifiers in AI outputs, allowing responsible tracking of content in cases with ethical or regulatory implications [286, 287, 33]. Also, proactive approaches that set standards and anticipates issues, as well as reactive approaches that address and rectify issues after they occur, can be utilized to achieve accountability [153].
@@
Regulations. AI regulation refers to laws, policies, and guidelines established to govern the development, deployment, and usage of AI. For example, European Union’s AI Act [244] categorizes AI applications by their risk levels, including unacceptable risk, high risk, general-purpose AI, limited risk, and minimal risk; the General Data Protection Regulation (GDPR) [243] enforces strict data protection and privacy requirements; and the Blueprint for an AI Bill of Rights [88] and NIST AI Risk Management Framework [154] provide guiding principles to foster responsible AI usage, emphasizing human oversight, transparency, and accountability.
@@
Access control. Access control in AI systems manages user access to data and models according to their permissions, thereby limiting unauthorized model access and data exposure. Effective methods include data access controls [101] and model access controls [6]. As an example, compositional fine-tuning enables each information silo (e.g., databases or documents) has its own fine-tuned model, allowing models to operate securely across multiple data silos and ensuring users access only to the fine-tunings they are authorized for.
@@
Interpretability, explainability, and transparency. Interpretability, explainability, and transparency are interrelated concepts but differ in their scopes. Interpretability focuses on understanding how AI models function internally and how they produce outputs [128, 74, 199, 185]. Explainability provides explanations to specific model outputs, presenting the results in ways that can be easily understood and trusted by end-users [39, 128, 74, 185, 273]. Transparency is the broadest concept that encompasses both interpretability and explainability and involves insights into the entire AI systems, such as how models make decisions, what data are used, and why specific results are produced [263, 114, 74, 185]. Transparency also extends to concepts like traceability and content provenance that track models and data to ensure accountability and reliability [219, 47, 49]. In practice, these concepts are often considered together to provide a more holistic understanding of AI systems [302, 59, 74, 185, 238]. Practices to improve transparency, interpretability, and explainability include a variety of methods, such as i) evaluating the contribution of various input elements (e.g., data features, user-defined concepts, or specific regions in images) to model outputs with methods such like removal-based explanations [43, 133, 128, 133, 206, 222, 48], counterfactual explanations [248, 102], concept activation explanations [107, 128], and saliency maps for visual data [3]; ii) X-of-Thought approaches that break down complex tasks into structured reasoning steps to make the reasoning process more interpretable to users, e.g., chain-of-thought [258, 295, 52, 68], tree-of-thought [131, 281, 28], graph-of-thought [17, 117], and their variants; iii) embedding watermarks to enhance traceability and content provenance for AI-generated outputs [286, 287, 47, 49]; iv) leveraging external knowledge sources, such as retrieval-augmented generation (RAG) through vector databases [98, 118], function calling [65, 125, 193], and web browsing [149], as reference for reasoning; v) leveraging interpretability algorithms or models as auxiliary models to generate insights into AI outputs and automate AI interpretability [200, 201, 20].
@@ 6 AI Safety in the 2020s: Challenges and Missions
This section utilizes the blueprint to guide our efforts today, navigating the challenges and opportunities of the current stage. As AI safety practitioners in the 2020s, we stand at a pivotal moment in history, with society moving toward an era where AI will be deeply integrated into daily life. While addressing current safety concerns, our primary mission is to anticipate and mitigate risks within the expanding AI landscape.
@@
A shift from AI safety to AI quality assurance. The rapid changing AI landscape requires a shift of our focus from narrowly defined AI safety concerns to a more comprehensive concept, i.e., AI quality assurance. Besides focusing on safety and security measures, practitioners can put more efforts on addressing alignment [161] of model outputs as well as the quality of interactions between users and AI to enhance the overall quality of AI services. Viewing AI safety as part of the quality assurance concept helps to enhance AI in a way that is safe, accurate, and capable of meeting the complex demands of real-world AI applications. The terminology “quality assurance” also guide us to think more about what we can do to align our current efforts with the blueprint picture.
@@
Alignment with more interdisciplinary collaborations. A key area that demands increased focus is the alignment of AI with deeper insights across different fields, people, and cultures. Current AI models, while advanced, often lack a deep understanding of ethical, social, and cultural contexts, while most of the alignment research tend to be conducted by people in the computer science field. However, more advanced AI systems necessitate interdisciplinary collaborations between AI researchers and experts from different fields such as psychology, sociology, history, art, and anthropology, etc, especially for AI applications for specific fields such as finance [268, 292], healthcare [236, 2], education [188, 139], etc. AI safety practitioners must work towards bridging this gap by fostering interdisciplinary collaborations. By expanding our focus beyond technical aspects, we can develop AI systems that are more aligned with human values as well as ethical and cultural dynamics of our society.
Hallucination. Hallucination remains a fundamental and persistent challenge, and is difficult to eliminate completely [92]. This is because hallucinations can be traced back to nearly every stage in the AI lifecycle, such as the quality and representativeness of training data, inappropriate user inputs, and the processes during inference. Low-quality or biased training data can cause models to generate incorrect information, while inappropriate user inputs may cause misleading details in the model outputs. During inference, models can also produce outputs that deviate from factual accuracy when they attempt to fill gaps in their training knowledge. Despite implementing Retrieval-Augmented Generation (RAG) [118] help mitigate hallucination, hallucination remains an inherent risk of AI models, making it crucial for content moderation strategies to identify and address these errors consistently at the current stage. Achieving this requires an ongoing effort on refining training data, grounding user queries properly, and implementing moderation to minimize the negative impact of hallucinations on the quality of the model outputs.
@@
Customization. Customization of AI workflow is essential for ensuring AI systems to meet the diverse user needs while operating effectively across different contexts and aligning with specific cultural norms, legal regulations, and ethical considerations. Current AI systems such as ChatGPT [137] tend to provide identical services across all user groups, regardless of individual differences. However, as AI becomes more integrated into everyday life, one-size-fits-all solutions are no longer adequate. People have different backgrounds, preferences, and requirements, and AI systems should be able to adapt to individual differences and provide safe, effective, and customized services. As an example, AI models interacting with children should respond differently compared with interacting with adults, which requires considering user age, learning objectives, cognitive abilities, emotional maturity, etc. In fields such as finance [268, 292], healthcare [236, 2], and education [188, 139], customizing AI services based on user profiles is crucial for enhancing safety while improving user experiences.
@@
Interaction. From our long-term goal towards the blueprint future, the advanced society involves complex interactions between intelligent robots and human users (similar ideas are also mentioned in [11, 70]). This calls for a shift in our focus from static, one-shot analyses to dynamic, context-aware interactions between users and models to enhance alignment. Beyond prompt-based AI systems and services, we can develop more sophisticated, context-aware interfaces for human-AI interactions to support multi-round interactions that adapt to evolving conversational contexts. Also, real-time safety protocols are in need to adjust to changing contexts and user needs. It is also essential to develop methodologies for safety assessment that functions effectively in complex and rapidly evolving environments, such as those involving interactions between multiple AI agents and human users.

### v4

@@ 5.3 Privacy
Privacy leakage in AI models refers to unintended exposure of sensitive information that the models may have learned during training. Adversaries can exploit model weights or gradients to infer sensitive information in the training data, e.g., data reconstruction attacks, membership inference attacks, etc [363, 105, 353]. Additionally, AI models are often trained on massive datasets that may contain sensitive data, such as personally identifiable information (PII). As a result, these models might memorize specific details of individuals, which might be reproduced inadvertently during interactions with users. Malicious users can also craft adversarial prompts to extract sensitive information from AI models with jailbreaks [151, 42].
@@
Addressing privacy leakage involves methods such as differential privacy [102, 297, 358, 356], federated learning [364, 354, 93, 352, 205], and privacy-preserving cryptographic protocols (e.g., homomorphic encryption [144, 228, 121]) for training AI models, ensuring that sensitive data cannot be extracted from weights or model outputs. Researchers also leverage differential privacy in in-context learning and finetuning [344, 355, 280], or design privacy-preserving prompts for querying LLMs [101, 63], to maintain data confidentiality during interactions between users and AI models. Other practices for ensuring data privacy include conducting rigorous data filtering before training, monitoring model inputs and outputs (see moderation in §5.2), etc, to ensure that the model outputs would not reveal sensitive information [176, 146, 95, 111] .
@@
When reflecting on the practices of privacy, a crucial aspect is the balance between privacy and utility. While ensuring data privacy is essential, it often comes at the expense of functionality, utility, and user experiences. In practice, implementing privacy measures in AI applications may fail to prevent malicious users while inadvertently degrading the experiences of benign users. Moreover, with advancements in technologies, breaching privacy becomes easier, while ensuring robust privacy protections is growing more complex and expensive. This highlights a deeper issue: those intent on exploiting vulnerabilities of AI models can often find new methods to bypass safeguards, whereas being overly cautious may impact benign users negatively.
@@ 5.4 AI Security and Defense Methods
Defense methods for ensuring AI security typically involve the following strategies: (1) aligning models with human values and ethical norms through Supervised Fine-Tuning (SFT) [361] and Reinforcement Learning from Human Feedback (RLHF) [204] to ensure the models to follow user instructions safely and responsibly; (2) leveraging built-in safeguards for content generation, e.g., exploring the decoding stage [106, 62], implementing reward-based mechanisms [188, 241], and leveraging hidden states [32, 62], to prevent the generation of unsafe content; (3) constructing datasets with safety-enhancing features, e.g., secure instructions and adversarial samples, and training models with such data to enhance their robustness against malicious inputs [329, 78, 176, 11, 215]; (4) employing prompt engineering methods, combined with chain-of-thought reasoning [43] or multi-agent systems [150], to help models understand user inputs better and react appropriately [223, 322, 310, 350]; (5) training models and/or incorporating moderation pipelines to detect unsafe content and prevent generating inappropriate outputs (see §5.2); (6) employing certified robustness approaches that provide formal guarantees on the model’s robustness against adversarial perturbations [87, 47]; and (7) utilizing differential privacy [1, 115, 267], homomorphic encryption [276, 121], and secure MPC [52, 134], to ensure that sensitive data are protected during model training, etc.
@@ 5.5 AI Serving Security
AI serving security involves secure deployment, operation, and maintenance of AI systems when providing services to users [209, 323]. It focuses on protecting models and interactions between users and models from various threats, such as supply chain vulnerabilities [211], model theft [208], model denial of service [207], and insecure plugins [206].
Supply chain vulnerabilities exist in software components, models, and training data provided by third party providers or user prompts collected with the supply chain [211]. Such vulnerabilities include using outdated models, deprecated third-party packages, improper handling of training data, etc, leading to data breaches, malicious injections, or even system failures [211]. Hackers may inject poisoned data through supply chain, introducing backdoors and biases during pretraining or fine-tuning [212]. Malicious users may craft prompts to bypass system controls, inducing unauthorized actions or unsafe model outputs [210].
@@
Distributed denial of service (DDoS) attacks compromise AI services, especially when models are integrated into widely accessible applications [207]. Attackers can craft prompts that exploit recursive behaviors in AI models, causing excessive computational resource consumption [244]. For example, an abnormal traffic pattern of a DDoS attack led to sporadic outages for OpenAI, affecting ChatGPT and developer tools for hours in November 2023 [213].
Plugins enhance model functionality by enabling interactions with external software, databases, web tools, or APIs [206]. However, they may introduce vulnerabilities when execution controls are inadequate. Adversaries may exploit insecure plugins with adversarial prompts to perform unauthorized actions, such as data exfiltration, remote code execution, and privilege escalation [206, 245]. Furthermore, without isolated environments, unauthorized access to plugins might even allow modifying system-level resources, which, in the worst cases, might cause the serving system to crash [245, 246].
@@ 5.6 AI Governance
AI governance is a framework of rules, practices, policies, and tools that ensure AI systems are built, developed, and used in a safe and responsible manner, aligning with social values and fulfilling legal standards [175, 24]. [272] defines 13 AI governance components in a continuous loop in AI life-cycle, including: 1) establishing accountability for AI; 2) assessing regulatory risks; 3) gathering inventory of use cases; 4) increasing values of underlying data; 5) assessing fairness and accessibility; 6) improving reliability and safety; 7) heightening transparency and explainability; 8) implementing accountability with human-in-the-loop; 9) supporting privacy and retention; 10) improving security; 11) implementing AI model lifecycle and registry; 12) managing risk; and 13) realizing AI value.
@@
Among these directions, items 1) and 8) are related to accountability; items 2), 6), 9), 10), and 12) are related to red teaming (see §5.1), content moderation and customized moderation pipeline (see §5.2), AI privacy (see §5.3), AI security and defense methods (see §5.4), and AI serving security (see §5.5); item 3) is related to business and is out of the scope of this subsection (see [272] for details on practical AI governance use cases); item 4) is related to data management such as training data processing [176], access control [6, 19, 203, 260], and data regulations [17, 303]; item 5) is related to moderation (see §5.2) and access control [6, 124]; item 7) is related to AI transparency; and items 11) and 13) are related to customizing AI workflow to enhance functionality and safety (see §5.2). Without loss of generality, this subsection discusses topics that are not mentioned in the previous subsections, including i) accountability, ii) regulations, iii) access control, and iv) transparency, interpretability and explainability.
@@
Accountability. Accountability refers to being responsible for actions and impacts of AI systems on individuals and society [238, 181]. Accountability can be defined as a relation of answerability with three conditions: authority recognition, interrogation, and limitation of power [196]. The accountability framework can be summarized with seven features (context, range, agent, forum, standards, process, and implications) and four key goals (compliance, reporting, oversight, and enforcement) [196]. These goals are often complementary, while policy-makers tend to focus on some goals over others, depending on specific objectives of AI governance.
@@
Accountability can be explored at different levels within the AI lifecycle, including data, model, and developers [237]. Typical approaches for enhancing accountability include algorithmic assessments, auditing, and data provenance techniques [237]. Proactive approaches and reactive approaches can be utilized to set standards and addressing issues after they occur [196]. Watermarking methods further support accountability by embedding traceable identifiers in AI outputs, allowing responsible tracking of content in cases with ethical or regulatory implications [346, 347, 36].
@@
Regulations. AI regulation refers to laws, policies, and guidelines established to govern development, deployment, and usage of AI. For example, European Union’s AI Act [299] categorizes AI applications by their risk levels, including unacceptable risk, high risk, general-purpose AI, limited risk, and minimal risk; the General Data Protection Regulation (GDPR) [298] enforces strict data protection and privacy requirements; and the Blueprint for an AI Bill of Rights [103] and NIST AI Risk Management Framework [197] provide guiding principles to foster responsible AI usage, emphasizing human oversight, transparency, and accountability.
@@
Access control. Access control in AI systems manages user access to data and models according to their permissions, thereby limiting unauthorized model access and data exposure. Effective methods include data access controls [124] and model access controls [6]. As an example, compositional fine-tuning enables each information silo (e.g., databases or documents) has its own fine-tuned model, allowing models to operate securely across multiple data silos and ensuring users access only to the fine-tunings they are authorized for [6].
@@
Interpretability, explainability, and transparency. Interpretability, explainability, and transparency are interrelated concepts but differ in their scopes. Interpretability focuses on understanding how AI models function internally and how they produce outputs [164, 83, 249, 229]. Explainability provides explanations to specific model outputs, presenting the results in ways that can be easily understood and trusted by end-users [44, 164, 83, 229, 331]. Transparency is the broadest concept that encompasses both interpretability and explainability and involves insights into the entire AI systems, such as how models make decisions, what data are used, and why specific results are produced [320, 142, 83, 229]. Transparency also extends to concepts like traceability and content provenance that track models and data to ensure accountability and reliability [272, 53, 55]. In practice, these concepts are often considered together to provide a more holistic understanding of AI systems [366, 68, 83, 229, 293].
Practices for enhancing transparency include a variety of methods, such as i) evaluating the contribution of input elements (e.g., data features, user-defined concepts, or specific regions in images) to model outputs with methods such like removal-based explanations [48, 170, 164, 170, 257, 275, 54], counterfactual explanations [304, 125], concept activation explanations [131, 164], and saliency maps (for visual data [3]); ii) X-of-Thought approaches that break down complex tasks into structured reasoning steps to make the reasoning process more interpretable to users, e.g., chain-of-thought [315, 357, 58, 77], tree-of-thought [167, 340, 31], graph-of-thought [20, 147], and their variants; iii) embedding watermarks to enhance traceability and content provenance for AI-generated outputs [346, 347, 53, 55]; iv) leveraging external knowledge sources, tools, or methods as reference for reasoning, e.g., retrieval-augmented generation (RAG) [117, 148], function calling [74, 157, 242], and web browsing [191]; v) leveraging interpretability algorithms or models as auxiliary tools to generate insights into AI outputs while automating AI interpretability [250, 251, 23].
@@ 6 AI Safety in the 2020s: Challenges and Missions
This section utilizes the blueprint to guide our efforts at the current stage. As AI safety practitioners in the 2020s, we stand at a pivotal moment in history, with society moving toward an era where AI will be deeply integrated into daily life. While addressing immediate safety concerns, our primary mission is to anticipate and mitigate risks within the expanding AI landscape. Thus, certain directions may require attention.
@@
A shift from AI safety to AI quality assurance. The rapid changing AI landscape requires a shift of our focus from narrowly defined AI safety concerns to a more comprehensive concept, i.e., AI quality assurance. Besides focusing on safety and security, practitioners can put more efforts on addressing alignment [204] as well as the quality of interactions between users and AI. Viewing AI safety as part of AI quality assurance helps enhance AI in a way that is safe, accurate, and capable of meeting the complex requirements of real-world AI applications. The terminology “AI quality assurance” also guides us to think more about what we can do to align our current efforts with the blueprint picture.
@@
Enhancing alignment with more interdisciplinary insights. Aligning AI with deeper insights across different fields, people, and cultures is important for enhancing the services provided by AI models. Current AI models, while advanced, often lack a deep understanding of ethical, social, and cultural contexts, while most of the existing work on alignment tend to be conducted by people in the computer science field. However, more advanced AI systems necessitate interdisciplinary collaborations between AI researchers and experts from diverse fields such as psychology, sociology, history, art, and anthropology, especially for AI applications that involve frequent interactions with different user groups, such as domains like finance [326, 354], healthcare [290, 2], education [235, 178], etc. AI safety practitioners might expand their focus beyond technical aspects and foster interdisciplinary collaborations, such that AI systems can be technically robust while being socially and ethically aligned with diverse user backgrounds.
Hallucination. Hallucination remains a fundamental and persistent challenge and is difficult to eliminate entirely, regardless of advancements in AI [107]. This is because hallucinations can originate at nearly every stage of AI lifecycle. Factors such as the quality and representativeness of training data, inappropriate user inputs, and inference processes all contribute to the occurrence hallucinations. Despite implementing RAG [148] helps mitigate hallucination, it remains an inherent risk of AI models. Thus, implementing a comprehensive moderation pipeline for identifying and addressing hallucinations is crucial, which involves grounding user queries with RAG, detecting presence of hallucinations, and fixing hallucinations in the model outputs if possible.
@@
Customization. Customizing AI workflows is essential to ensure that AI systems address diverse user needs effectively and align with specific cultural norms, legal regulations, and ethical considerations. Current AI systems such as ChatGPT [176] tend to provide similar or identical services to all user groups, regardless of individual differences. However, as AI becomes increasingly integrated into daily life, one-size-fits-all solutions are inadequate. With diverse backgrounds and needs, users might expect AI systems to adapt to their individual differences and preferences, and deliver safe, effective, and personalized services. As an example, AI models interacting with children should respond differently compared with interacting with adults, necessitating considering user age, learning objectives, cognitive abilities, emotional maturity, etc. In domains such like finance [326, 354], healthcare [290, 2], and education [235, 178], customizing AI services based on user profiles is crucial for enhancing safety while improving user experiences.
@@
Interaction. Our long-term vision for the blueprint future suggests that the Internet of Everything involves complex interactions between intelligent robots and human users (similar ideas also discussed in [12, 79]), which calls for a shift in our focus from static, one-shot analyses to dynamic, context-aware interactions between users and models to improve alignment. Beyond prompt-based AI systems and services, we can develop more sophisticated, context-aware interfaces for human-AI interactions to support multi-round interactions that adapt to evolving conversational contexts. Also, real-time safety protocols are in need to adjust to changing contexts and user needs. It is also essential to develop methodologies for safety assessment that functions effectively in complex and rapidly evolving settings, such as those involving interactions between multiple AI agents and human users.